# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

### v1

@@ Abstract
We study the Ergodic Properties of Random Walks in stationary ergodic environments without uniform ellipticity under a minimal assumption. There are two main components in our work. The first step is to adopt the arguments of Lawler to first prove a uniqueness principle. We use a more general definition of environments using Environment Functions. As a corollary, we can deduce an invariance principle under these assumptions for balanced environments under some assumptions. We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk. The second is to transfer the results we deduce in balanced environments to general ergodic environments(under some assumptions) using a control technique to derive a measure under which the local process is stationary and ergodic. As a consequence of our results, we deduce the Law of Large Numbers for the Random Walk and an Invariance Principle under our assumptions.
@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments(RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk, and the second question is about the probability of the walk escaping to infinity along a given direction. Both have been studied in detail in $d=1$ by Solomon[11] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
Varadhan in [12] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$. Furthermore, Nicholas Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0. These results assume that the environment is uniformly elliptic and i.i.d.. In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
@@
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This section is one of the most fundamental contributions, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
@@
Our definition of environments allows us to transfer our results from the balanced case to the general environment, by coupling the random walk and a reflected walk. We then combine a novel technique which we shall name as the Action of Linear Transformations and another control technique to execute a proof.
@@ 2. Prerequisites
$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,...,d.$
@@
Therefore, the generators of this Markov Chain is given by
$L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)(g(x+v)-g(x)),$
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},\omega)(x)$ as follows
$c(\mathscr{E},\omega)(x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
@@ 3. Main Results and Discussions
These results notably weaken the assumptions of ellipticity and derives an alternate proof of Theorem 1 of [5] for $d\geq 2$, which states that
@@
To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transcience of balanced elliptic walks. The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
@@
To state the next result, which connects balanced and general unbalanced regimes, we need the notion of embedded environments. The motivation for the construction can be understood only after the reader gets acquainted with the control argument which will be demonstrated accordingly.
@@ 3.1. Embedded Environments
    $\lim_{n\to\infty}\frac{X_{\omega}(n)}{n}=v\quad\text{$\omega-\mu^{0}$-a.e}.$
@@ 3.2. Remarks:
The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$ namely the original and a reflected walk. This is motivated by the fact that the averaged random walk is balanced. Through the control argument, we pass on from the balanced to the general case. The Action of Linear Transformations proves crucial in performing this step. This might and should appear strange without prior context, but we shall see this arising naturally as an artifact of our proof technique.
@@ 4. An outline of the Proofs
The proof of Theorem 3.5 is postponed and will be shown as a natural consequence of our uniqueness principle. Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7. We divide the proof into sections. Although each section may be read independently without loss of continuity, we recommend following the order presented, as it provides the intended motivation.
@@
The two main themes in this proof are the use of Periodised Environments as used in [7] and our uniqueness principle. We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof. The proof in [7] adopts the arguments of Papanicolaou and Varadhan [8] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments. Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle. The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampere Equation by Lawler. We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3. We further use this uniqueness principle to derive Theorem 3.5. We emphasise that Theorem 3.5 is not a result on Random Walk in Random Environments but on general balanced elliptic walks.
@@
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. The control argument allows us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
@@ 5.1. Uniqueness Principle:
Recall that $Y_{j}$ is a Markov chain with the generator
@@ 5.2. Periodised Environments
$(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,..d.$
@@
The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the generator $L_{\mathscr{E}(\hat{\omega})}$.
@@
First we see the proof of 5.3, assuming $5.$
@@
By log-convexity of $l_{p}$ norms and Holder’s Inequality,
@@
We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section
@@
To see the equivalence one notes the fact that $z(x)=Qf$.
Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampere Operator for a given concave function $z$ given by
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\penalty 10000\ ||R_{n}g||_{\infty}.$
The last line follows from the Strong Markov property as
@@
$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right)$
@@
Further more observe that since $x\in T_{n}$ we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
@@
By construction, the assumptions of Lemma 5.5 are satisfied, thereby the proof.
Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a proof of Theorem 3.5.
@@
Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed consider
@@
It is again easy to show that
@@
Denote by the $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one where it assigns 0.
Therfore we have that
@@
Denote by $\{X_{j}\}_{j\geq 0}$ to be the simple symmetric random walk. Then we have by Theorem 5.1 that
@@
Similarly we note that
@@
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}f(X_{j})]$
@@
We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof we shall assume $\mathscr{E}=\operatorname{Id}$.
@@ 8.1. Action of Linear Transformations:
The induced random walk under $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
@@
Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{n})}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
@@
$\{\tau_{e}{T.\omega}\in T(B)\}=\{T(\tau_{Tx}\omega)\in T(B)\}=\{\tau_{Tx}(B)\in B\}\quad\penalty 10000\ \forall e\in V_{d}.$
@@
If $\tau_{X_{\mathscr{E}(\omega)}(n)}\omega$ has a stationary measure $Q$, $\tau_{T(X_{\mathscr{E}(\omega)}(n))}\omega$ is also stationary under $Q\circ T$.
@@
But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.) Now we note that from the martingale property and the Markov property,
@@
Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 1}$. The first make an observation on this measure.
Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,...,d$.
@@
This shows $\lambda^{\prime}\circ\alpha$ is also an invariant stationary measure for the environment process. Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$. One can also show that $\mu\circ\alpha$ is stationary ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
@@
We redefine $X_{\Lambda(\omega)}(n)$ in the following manner.
@@
where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ are the generators of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk. Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d. Bernoulli random variables.
@@
$K_{d}^{\Lambda}(\omega,B)\Big|\gamma=K_{\gamma}(\omega,B)=\gamma K_{d}(\omega,B)+(1-\gamma)G_{d}(\omega,B).$
@@
Let $n\geq 1$. Let $B_{0},B_{1},...,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated to $\omega$ be $\lambda^{\prime}$. Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$. Then for every $1\leq k\leq n-1$, we have
@@
$\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]=\frac{1}{2^{n}}\sum_{(\gamma_{1},...,\gamma_{n})\in\{0,1\}^{n}}\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]$
The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$. Now conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$. It is clear that
$\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,....,n\}=\tau_{X_{\omega}(n)}\omega.$
Therefore it is now enough to establish Lemma 9.3.
@@
Suppose the result is true for $n=m$. Let $n=m+1$. Then we must verify the statement for $k=m+1$ only. For $k=m+1$, we write
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
Perform the substitution $(\omega_{m},...,\omega)\mapsto(\alpha.\omega_{m},...,\omega)$ to get
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
@@
$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},...,\gamma_{m}\Big]$
@@
Acknowledgements:
The author is deeply indebted to Dr. Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for helpful discussions and his immense help with the manuscript. The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.

### v2

@@ Abstract
We study the Ergodic Properties of Random Walks in stationary ergodic environments without uniform ellipticity under a minimal assumption. There are two main components in our work. The first step is to adopt the arguments of Lawler to first prove a uniqueness principle. We use a more general definition of environments using Environment Functions. As a corollary, we can deduce an invariance principle for balanced environments under the assumptions. We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk. The second is to transfer the results we deduce in balanced environments to general ergodic environments(under some assumptions) using a control technique to derive a measure under which the local process is stationary and ergodic. As a consequence of our results, we deduce the Law of Large Numbers for the Random Walk and an Invariance Principle under our assumptions.
@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments (RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk. Thid have been studied in detail in $d=1$ by Solomon[11] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
Varadhan in [12] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$. Furthermore, Noam Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0. These results assume that the environment is uniformly elliptic and i.i.d.. In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
@@
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This is one of the key contributions of this p, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
@@
Our definition of environments allows us to transfer our results from the balanced case to the general environment by coupling the random walk and a reflected walk. We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof.
@@
Acknowledgements:
The author is an undergraduate student at the Indian Statistical Insitute, Kolkata. This work was completed during the third-year of his studies. He is deeply indebted to Dr. Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript. The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.
@@ 2. Prerequisites
$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,\dots,d.$
@@
$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}$
@@
Therefore, the kernel of this Markov Chain are given by
$L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)g(x+v),$
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},x)$ as follows
$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
@@ 3. Main Results and Discussions
These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [5] for $d\geq 2$, which states that
@@
To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transience of balanced elliptic walks. The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
@@
To state the next result, which connects balanced and general unbalanced regimes, we need the notion of embedded environments. The motivation for the construction can be understood only after the reader gets acquainted with the control argument, which will be demonstrated accordingly.
@@ 3.1. Embedded Environments
    $\lim_{n\to\infty}\frac{X_{\omega}(n)}{n}=v\quad\omega\penalty 10000\ \text{$\mu$-a.e}.$
@@ 3.2. Remarks:
The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$, namely the original and a reflected walk. This is motivated by the fact that the averaged random walk is balanced. Through the control argument, we pass from the balanced to the general case. The Action of Linear Transformations proves crucial in performing this step. This might and should appear strange without prior context, but we shall see this arising naturally as an artifact of our proof technique.
@@ 4. An outline of the Proofs
The proof of Theorem 3.5 is postponed and will be shown as a natural consequence of our uniqueness principle. Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7. We divide the proof into various sections in order to properly motivate the proof.
@@
The two main themes in this proof are the use of Periodised Environments as used in [7] and our uniqueness principle. We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof. The proof in [7] adopts the arguments of Papanicolaou and Varadhan [8] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments. Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle. The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler. We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3. We further use this uniqueness principle to derive Theorem 3.5. We emphasise that Theorem 3.5 is not a result on Random Walk in Random Environments but on general balanced elliptic walks.
@@
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. This technique allows us to control the paths of the random walk which shall be key to the proof. The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
@@ 5.1. Uniqueness Principle:
Recall that $Y_{j}$ is a Markov chain with the kernel
@@ 5.2. Periodised Environments
$(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,\dots,d.$
@@
The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\hat{\omega})}$.
@@
First, we see the proof of 5.3, assuming $5.$
@@
By log-convexity of $l_{p}$ norms and Hölder’s Inequality,
@@
We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section.
@@
To see the equivalence, one notes the fact that $z(x)=Qf$.
Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampère Operator for a given concave function $z$ given by
@@
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}$
The last line follows from the Strong Markov property since
@@
$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
@@
Furthermore, observe that since $x\in T_{n}$, we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
@@
By construction, the assumptions of Lemma 5.5 are satisfied, thereby completing the proof.
Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a natural proof of Theorem 3.5.
@@
Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed, consider
@@
It is again easy to show that.
@@
Denote by $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one, where it assigns 0.
Therefore, we have that
@@
Denote by $\{X_{j}\}_{j\geq 0}$ the simple symmetric random walk. Then we have by Theorem 5.1 that
@@
Similarly, we note that
@@
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]$
@@
We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof and from now on we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.
@@ 8.1. Action of Linear Transformations:
The induced random walk under the transformation $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
@@
Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{\omega}(n))}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
@@
$\{\tau_{e}{T.\omega}\in T(B)\}=\{T(\tau_{T.e}\omega)\in T(B)\}=\{\tau_{T.e}\omega\in B\}\quad\penalty 10000\ \forall e\in V_{d}.$
@@
If $\tau_{X_{\omega}(n)}\omega$ has a stationary ergodic measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary and ergodic under $Q\circ T$.
@@
But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property and the fact that $\{h(\bar{\omega}(n)\}_{n\geq 0}$ is a martingale,
@@
Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. First, we make an observation on this measure.
Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,\dots,d$.
@@
This shows $\lambda^{\prime}\circ\alpha$ is also an invariant stationary measure for the environment process. Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$. One can also show that $\mu\circ\alpha$ is stationary and ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
@@
We redefine $\{X_{\Lambda(\omega)}(n)\}_{n\geq 0}$ in the following manner.
@@
where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ kernels of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk. Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d. Bernoulli random variables.
@@
$K_{d}^{\Lambda}(\omega,B)\Big|\gamma:=K_{\gamma}(\omega,B)=\gamma K_{d}(\omega,B)+(1-\gamma)G_{d}(\omega,B).$
@@
Let $n\geq 1$. Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated with $\omega$ be $\lambda^{\prime}$. Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$. Then for every $1\leq k\leq n-1$, we have
@@
$\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\gamma_{1},\dots,\gamma_{n}\right]=\frac{1}{2^{n}}\sum_{\{\gamma_{i}\}_{i=1}^{n}\in\{0,1\}^{n}}\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\{\gamma_{i}\}_{1\leq i\leq n}\right]$
The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$. Now, conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$. It is clear that
$\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,\dots,n\}=\tau_{X_{\omega}(n)}\omega.$
Therefore, it is now enough to establish Lemma 9.3.
@@
Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity as others follow from the technique of proof. For $k=m+1$, we write
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
Perform the substitution $(\omega_{m},\dots,\omega)\mapsto(\alpha.\omega_{m},\dots,\omega)$ to get
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
@@
$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},\dots,\gamma_{m}\Big]$